### v1

# Countering adversarial evasion in regression analysis

Abstract
Adversarial machine learning challenges the assumption that the underlying distribution remains consistent throughout the training and implementation of a prediction model. In particular, adversarial evasion considers scenarios where adversaries adapt their data to influence particular outcomes from established prediction models, such scenarios arise in applications such as spam email filtering, malware detection and fake-image generation, where security methods must be actively updated to keep up with the ever-improving generation of malicious data. Game theoretic models have been shown to be effective at modelling these scenarios and hence training resilient predictors against such adversaries. Recent advancements in the use of pessimistic bilevel optimsiation which remove assumptions about the convexity and uniqueness of the adversary’s optimal strategy have proved to be particularly effective at mitigating threats to classifiers due to its ability to capture the antagonistic nature of the adversary. However, this formulation has not yet been adapted to regression scenarios. This article serves to propose a pessimistic bilevel optimisation program for regression scenarios which makes no assumptions on the convexity or uniqueness of the adversary’s solutions.
1 Introduction
Adversarial machine learning considers the exploitable vulnerabilities of machine learning models and the strategies needed to counter or mitigate such threats [32]. By considering these vulnerabilities during the development stage of our machine learning models, we can work to build resilient methods [9, 11] such as protection from credit card fraud [35] or finding the optimal placement of air defence systems [20]. In particular, we consider the model’s sensitivity to changes in the distribution of the data. The way the adversary influences the distribution can fall under numerous categories, see [21] for a helpful taxonomy that categorises these attacks. We focus on the specific case of exploratory attacks, which consider the scenarios where adversaries attempt to modify their data to evade detection by a classifier. Such attacks might occur in security scenarios such as malware detection [3] and network intrusion traffic [31]. In a similar vein, and more recently, vulnerabilities in deep neural networks (DNN) are being discovered, particularly in the field of computer vision and image classification; small perturbations in the data can lead to incorrect classifications by the DNN [33, 19]. These vulnerabilities raise concerns about the robustness of the machine learning technology that is being adopted and, in some cases, in how safe relying on their predictions could be in high-risk scenarios such as autonomous driving [15] and medical diagnosis [16]. By modelling the adversary’s behaviour and anticipating these attacks, we can train classifiers that are resilient to such changes in the distribution before they occur.
Game theory provides an effective and popular technique to model adversarial scenarios. These games see one player, the learner, attempt to train a prediction model while another player, the adversary, modifies their data in an attempt to influence particular outcomes from the learner’s predictor. See, for example [10] for early work in modelling adversarial learning. The precise structure of such a model then depends on a number of factors, such as whether the attack occurs either at training time [27], or implementation time [5]. Further to this, while some games assume the players act simultaneously [4, 29], others allow for sequential play where either the adversary acts first [25, 26, 7, 24] or the learner acts first [21, 5]. We focus our attention on attacks made at implementation time where the adversary seeks to influence an already established prediction model, commonly known as evasion attacks in the literature [21]. Sequential evasion games naturally take on the form of a bilevel optimisation program. In this formulation, the learner, taking the role of the leader in the upper-level, train a prediction function while anticipating how the adversary, taking the role of the follower in the lower-level, will react. The solution to this program then results in a resilient prediction function. Pessimistic bilevel optimisation in particular has proved particularly effective. When multiple optimal solutions are available to the adversary, the pessimistic model, unlike its optimistic variant, assumes no cooperation between the adversary and the learner. This formulation better captures the antagonistic nature of adversarial evasion scenarios [2, 5].
In this article, we adapt the pessimistic bilevel program proposed by [37] to train resilient predictors in regression scenarios. We model scenarios where we might expect adversaries to modify their data in an attempt to influence a particular label from a prediction model. We use the motivating example of an adversary providing falsified information about a house in order to influence a higher valuation. We explore properties of the constrained bilevel model under new choices of objective functions, showing the possible existence of multiple optimal solutions and hence the benefit of a constrained pessimistic bilevel approach, which, to the best of our knowledge, has not yet been applied to regression analysis. Moreover, we take this opportunity to exploit the fact that our method does not require the use of feature maps in its solution method. Feature maps, such as PCA, as was used [5], are typically not bijective. Consequently, although a solution to the adversary can be found in the map-induced feature space, a corresponding solution in the original feature space cannot be identified. However, under our model, we can view the explicit values of the adversary’s data at their solution. These values can give us insight into the strategies played by the adversary and allows us to identify the extent to which each feature needs to be modified in order to adequately influence the prediction model. From this, we can identify which features are most vulnerable. For example, if a certain feature required only a minor perturbation, then we might consider this feature to be particularly vulnerable to adversarial influence. On the other hand, if a feature required large transformations to trick the learner, then we might consider this feature to be resilient to attacks.
The contributions of this article can be summarised as follows. We adapt the bilevel model proposed in [37] to accommodate regression analysis and present a novel pessimistic bilevel optimisation program to train prediction functions that are resilient to strategic changes in the data of regression scenarios. This bilevel program makes no assumptions on the convexity of the lower-level problem, or the uniqueness of its solution, which, to the best of our knowledge, has not previously been proposed for adversarial regression analysis. We design experiments to assess the performance of our model and showcase its ability to outperform existing methods. Finally, we demonstrate a key benefit of our model over existing approaches. Where existing approaches required the use of kernel transformations to ensure convexity, we are able to keep the lower-level variable (the adversary’s data) in its original feature space, allowing us to investigate the value of the data at the solution. From this we can infer which features require the greatest transformation in order to effectively influence the learner’s prediction function.
The remainder of this paper is outlined as follows; we begin in Section 2 by presenting the constrained pessimistic bilevel model to train resilient prediction functions in regression analysis before outlining the solution method. Following this, we assess the performance of the bilevel model on two datasets in Section 3 before concluding our findings in Section 4.
2 Training resilient prediction functions
In this section we present an adaptation of the constrained pessimistic bilevel program proposed in [37] to model adversarial regression. In particular, we demonstrate how the process of adversarial training can be used to train resilient predictors under the scenario of adversaries attempting to strategically modify their data in an attempt to be assigned incorrect labels from established prediction models. Such scenarios might arise, for example, in property valuation. An adversary might lie and provide falsified information in an attempt to influence a higher valuation. Through suitable choices of objective functions, we construct a constrained pessimistic bilevel program that sees a learner, in the upper-level, train a prediction function while considering how an adversary, in the lower-level, might alter their data towards a target outcome from the prediction model. The adversary is assigned some existing data as a starting point. The constraints are then used to measure the similarity between the adversary’s solution and starting point. As demonstrated in [37], these constraints ensure that the adversary’s data remain plausibly realistic and prevent instances of data losing their intended meaning. In this way, the bilevel model more realistically simulates adversarial transformations, leading to improved performance. To the best of our knowledge, a constrained pessimistic bilevel program with no assumptions on convexity or uniqueness of lower-level solution has previously been proposed for adversarial evasion scenarios in regression analysis.
Let $x\in\mathbb{R}^{q}$ be a sample (row) of data containing $p\in\mathbb{N}$ features with corresponding label $y\in\mathbb{R}$. For some weights $w\in\mathbb{R}^{q}$, we define the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, as the linear weighted sum of the features,
$\sigma(w,x):=w_{0}+w^{T}x.$ (1)
This prediction function is a common choice, for example, in linear regression. We then construct a loss function, $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, to train the prediction function and identify appropriate weights. As in linear regression, we can achieve this by minimising the squared error between the learner’s prediction of $x$ and the label, $y$,
$\mathcal{L}(\sigma(w,x),y):=(\sigma(w,x)-y)^{2}=(w_{0}+w^{T}x-y)^{2}.$
The adversary, in the lower-level, transforms some data with the aim of having the learner’s prediction function, $\sigma$ output a desired target label. For example, consider again the example of an adversary lying about features of a housing property to influence a higher valuation than deserved. If, for example, the actual valuation price is $y\in\mathbb{R}$, the adversary might set their target, $z\in\mathbb{R}$, as $z=y+\nu$, where $\nu>0$. The lower-level (adversary’s) loss function is then defined as the squared distance between the learner’s prediction and the target label,
$\ell(\sigma(w,x),z):=(\sigma(w,x)-z)^{2}=(w^{T}x-z)^{2}.$ (2)
With the loss functions established, we now work them into objective functions that accommodate a full dataset. We divide the training data into two sets, a static set as would typically be used in the training of a machine learning predictor, and then second set of data which can be manipulated by the adversary. Let $D\in\mathbb{R}^{nq},D=(D_{1},\dots,D_{n})^{T}$ be the static set of $n\in\mathbb{N}$ instances of data where each $D_{i}\in\mathbb{R}^{q},\;i=1,\dots,n$, is a row vector containing the values of $q\in\mathbb{N}$ features, and let $\gamma\in\{0,1\}^{n}$ be the corresponding corresponding collection of labels. Let $X\in\mathbb{R}^{mq}$ be the set of $m\in\mathbb{N}$ instances which can be manipulated by the adversary, defined as,
$X:=\begin{pmatrix}X_{1}^{T}\\ \vdots\\ X_{m}^{T}\end{pmatrix}$,
where each $X_{i},\;i=1,\dots,m$, is a row vector of features. Let $Y\in\{0,1\}^{m}$ be the corresponding labels of the adversary’s data. The upper level objective of the bilevel program sees the learner minimising their loss over both sets of data. Note that the learner considers the true labels of the adversary’s data, where $F:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ is the upper-level objective,
$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i}).$
Then, given a set of weights, $w$ of the learner’s prediction function, the adversary, in the lower-level seeks to modify their data such that the learner’s prediction moves towards their target labels, $Z$. Let $f:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ be the adversary’s objective, which is defined as the sum of the adversary’s loss over their data,
$f(w,X):=\frac{1}{m}\sum_{i=1}^{m}\ell(\sigma(w,X_{i}),Z_{i}).$ (3)
Finally, we define the constraints $g:\mathbb{R}^{mq}\rightarrow\mathbb{R}^{m}$ as
$g(X):=\begin{pmatrix}g_{1}(X_{1})\\ \vdots\\ g_{m}(X_{m})\end{pmatrix}$, (4)
where each constraint function, $g_{i}:\mathbb{R}^{q}\rightarrow\mathbb{R},\;i=1,\dots,m$, is defined as
$g(X):=\delta-d(X_{i},X^{0}_{i}),$ (5)
where $d:\mathbb{R}^{q}\times\mathbb{R}^{q}\rightarrow\mathbb{R}$ is some similarity function. These constraints serve to measure and restrict the cosine similarity between the adversary’s solution and initial value to be below some threshold $\delta\in\mathbb{R}$. As outlined in [37], these constraints prevent the adversary’s data becoming nonsensical and too far removed from its intended message. This continues to be useful under the regression scenarios as we can reduce the likelihood of the adversary producing unrealistic data. For example, in reality, an adversary cannot claim that their house was built thousands of years ago when in reality it was built a mere 5 years ago. This would be immediately suspicious to anybody who checks the adversary’s data. However, smaller transformations, such as within the range of $\pm 2$ years would be far more believable. Similarly to the work in [37], we use the cosine similarity to measure the extent to which the adversary has modified their data. Let $X^{0}\in\mathbb{R}^{mq}$ be the initial value of the adversary’s data, then the constraints are then defined as follows,
$g_{i}\left(X\right)=\delta-d(X_{i},X^{0}_{i})=\delta-\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i\in\{1,\dots,m\}.$ (6)
Then, setting the initial value of the adversary’s data to some existing data in the training set, we can be confident that under these constraints, the adversary produces data that are similar in nature to some real example.
We now formulate the interactions between the learner and adversary as a bilevel optimisation problem. To begin, given a set of weights, $w$, of the learner’s prediction function, let $S(w)$ be the set of optimal data for the adversary,
$S(w):=\operatorname*{argmin}_{X\in\mathbb{R}^{mq}}\left\{f(w,X)\mid g(X_{i})\leq 0\right\},$ (7)
This set contains all instances of data that will trick the learner into assigning the adversary their desired labels $Z$. For example, the falsified falsified report that gives a property a better valuation than it actually deserves. Note that the adversary’s objective function under the linear regression loss in (2) and prediction function (1) possess multiple optimal solutions. This can be directly shown using in [37, Proposition 1] since the prediction function, $\sigma$, can be expressed as a function of the linear combination $w_{0}+w^{T}x$ for some $x\in\mathbb{R}^{q}$. Simply let $\Omega(w_{0}+w^{T}x)=w_{0}+w^{T}x$ and the proof follows.
Note also that its has already demonstrated in [37, Proposition 2] that the feasible region defined by these constraints is non-convex. Therefore, in the case of multiple optimal solutions to the adversary’s problem, given by the set $S(w)$, we make the pessimistic assumption that the adversary will not act cooperatively, namely, they will choose the solution which maximises the learner’s objective, while the learner seeks to minimise it. The complete bilevel program is then given by
$\min_{w\in\mathbb{R}^{p}}\max_{X\in S(w)}F(w,X).$ (8)
The non-convexity and existence of multiple optimal solutions to the lower-level problem motives the use of the solution method proposed in [37]. To begin, we collect the upper and lower variables, $w$ and $X$, into the block variable $z$ as below
$z:=\begin{bmatrix}w\\ X\end{bmatrix}\in\mathbb{R}^{q+mq},\qquad$
and define the block variable $\zeta\in\mathbb{R}^{2b+1}$ as,
$\xi:=\begin{bmatrix}\beta\\ \hat{\beta}\\ \lambda\end{bmatrix}\in\mathbb{R}^{2b+1},$
where $\beta,\hat{\beta}\in\mathbb{R}^{b}$ and $\lambda\in\mathbb{R}$ are lagrange multiples with associated Lagrangian-type functions $L^{\textup{p}}_{w},L^{\textup{p}}_{X},\ell^{\textup{p}}\colon(\mathbb{R}^{q}\times\mathbb{R}^{mq})\times\mathbb{R}^{2b+1}\to\mathbb{R}$ as stated below:
$L^{\textup{p}}_{w}(z,\xi):=-F(w,X)+(\beta-\lambda\hat{\beta})^{\top}g(X)$
$L^{\textup{p}}_{X}(z,\xi):=-F(w,X)+\lambda f(w,X)+\beta^{\top}g(X)$
$\ell^{\textup{p}}(z,\xi):=f(w,X)+\hat{\beta}^{\top}g(X).$
Now, we can define the block functions, $\mathcal{G}$ and $\mathcal{H}$ as below,
$\mathcal{G}(z):=\begin{bmatrix}g(X)\\ g(X)\\ 0\end{bmatrix},\qquad\mathcal{H}(z,\xi;B_{i}):=\begin{bmatrix}\nabla_{w}L^{\textup{p}}_{w}(z,\xi)\\ \nabla_{X}L^{\textup{p}}_{X}(z,\xi)\\ \nabla_{X}\ell^{\textup{p}}(z,\xi)\end{bmatrix}$.
Then, as outlined in the solution method in [37, Section 3], we can find stationary points of the bilevel program by solving the following system,
$\zeta\geq 0,\,\mathcal{G}(z,\zeta)\leq 0,\,\zeta^{T}\mathcal{G}(z,\zeta)=0,\,\mathcal{H}(z,\zeta;B_{i})=0.$
This system can be converted into a system of equalities via the Fischer-Burmeister function,
$\Phi(z,\zeta;B_{i})=\begin{pmatrix}\mathcal{H}(z,\zeta;B_{i})\\ \varphi_{\textup{FB}}\left(\zeta,-\mathcal{G}(z,\zeta)\right)\end{pmatrix}=0,$ (9)
where $\xi:=(z,\zeta)\in\mathbb{R}^{p}\times\mathbb{R}^{m\times p}\times\mathbb{R}^{2b+1}$ is a block variable and $\varphi_{\textup{FB}}$ is the Fischer-Burmeister, defined as
$\varphi_{\textup{FB}}(a,b)=\sqrt{a^{2}+b^{2}}-a-b.$
This system can then be solved using the Levenberg-Marquardt algorithm given in [37, Algorithm 1], where $\Psi:(\mathbb{R}^{p}\times\mathbb{R}^{mp})\times\mathbb{R}^{2b+1}$ is defined as,
[…]
The resulting weights of this solution can be used to construct a prediction function that has accounted for adversarial influence during the training process. With the inclusion of the constraints in (6), we can be confident that the adversary’s data retains the intended nature and realism after the transformation.
3 Numerical experiments
We assess the resilience of the bilevel model to strategic adversarial modifications. To do so, we design experiments on two datasets, the first of which, named Wine Quality records 11 features of 4898 bottles of red wine such as acidity and alcohol content. A corresponding quality score between 0100-10 is measured and recorded for each bottle. For this dataset, we consider scenarios where wine producers might lie when reporting the features of the wine or perhaps bribe quality assessors to alter the results in order to appear to be producing wine of a higher quality than they actually are. The second dataset, named Real Estate, contains the values of 6 features of 414 houses such as their age and distance to the nearest train station along with the corresponding sale price. We consider scenarios where sellers might lie about features of the house in order to be able to charge a higher price than deserved.
Further to this, we use the numerical experiments in this section as an opportunity to explore a particularly useful aspect of the constrained model which was not exploitable in [37]. Specifically, we demonstrate how, unlike existing pessimistic bilevel approaches to adversarial training, our models allow us to investigate which features of the adversary’s data were changed and hence give us insight into the most vulnerable aspects of the prediction models. In the current proposal for a pessimistic bilevel approach to evasion attacks, [5], pass the adversary’s data through a feature map. This plays a vital role in the proposed solution process, as the resulting feature space can be assumed to be unrestricted, which is convex. This, coupled with a strictly convex objective function, grantees a unique solution to the adversary’s problem in the lower-level, allowing the model to be reorganised as a pessimistic bilevel optimisation problem, which is easier to solve. However, since our model makes no assumptions on the convexity of the adversary’s problem or the uniqueness of its solution, we can leave the adversary’s data in its original feature space. Therefore, we can observe the adversary’s solution and infer the extent to which features must be changed in order to trick the learner. This analysis could not be performed in the experiments in [37], as the text-based data needed to be embedded before analysis. This embedding is not bijective and, consequently, the value of the adversary’s data at the solution provided no useful insight into the adversary’s strategy. However, in this chapter, we take the opportunity to test the bilevel model on some datasets that require no feature manipulation and hence allow us to investigate the adversary’s movements.
We divide the data into a training, and test set by the ration of $80\%$ and $20\%$ respectively and simulate adversarial attacks to inject the into the test set. Let $X^{\text{test}}\in\mathbb{R}^{T\times p}$ be the test set containing $T$ samples. Before evaluation, we simulate adversarial influence on the test set by transforming a portion of the instances towards target values $Z^{\text{test}}=Y^{\text{test}}+\Delta$, defined as a perturbation of the ground truth test labels $Y^{\text{test}}$, where $\Delta\in\mathbb{R}$ is a perturbation. Let $t<=T$ be the portion of data modified by an adversary and let $I\subset\{1,\dots,T\}$ be the corresponding indexes of the test data which are manipulated. We simulate a range of adversarial abilities. We generate adversarial test data by modifying $X^{\text{test}}_{I}$ towards the adversary’s target labels. To simulate a range of abilities, we randomly generate each adversary their own similarity threshold from the range $(0.8,1)$. Let $\delta^{\text{test}}\in(-1,1)^{t}$ be the set of similarity thresholds where $\delta_{i}\sim U(0.8,1)\;\forall\;i\in I$ and $U$ denotes the uniform distribution. The instances of adversarial test data and then found by solving the problem
$X^{\text{test}}_{i}:=\operatorname*{argmin}_{x\in\mathbb{R}^{q}}\left\{\ell(\sigma(w,x),Z^{\text{test}}_{i})\,\mid\,g(x,x^{0};\delta^{\text{test}}_{i})\leq 0\right\},\quad i\in I,$
where $x^{0}\in\mathbb{R}^{q}$ is the initial values of $x$. We set $T$ to be $10\%$ of the size of the test set. All features and labels are normalised to the range $(0,1)$ and we set the adversarial perturbation to be $\Delta=2\text{std}(Y^{\text{test}})$ where $\text{std}(Y^{\text{test}})$ is the standard deviation of the test labels. We compare our model to that of a typically trained linear regression predictor, named LinReg, for various values of $m$ and plot in Figure 1. We can see for both datasets a similar trend of an initially high mean square error (MSE) for $m=1$, before a drastic decrease to an optimal value of $m=2$. Following this, we see a gradual increase in MSE as $m$ increases, before plateauing around an MSE roughly equal to that of the traditional predictor.
Figure 1: Comparison of the pessimistic bilevel model with existing solvers
We now investigate the movement of each feature by the adversary during the training process to gain insight into which features are most vulnerable. For each feature (column), we calculate the average absolute distance between the adversary’s solution $X^{*}$ and the original position $X^{0}$,
$\frac{1}{m}\sum_{j=1}^{m}\left|X^{*}_{ij}-X^{0}_{ij}\right|,\quad i=1,\dots,m.$
Under this measure, a larger distance implies that the adversary needed to transform the data by a larger extent in order to sufficiently fool the leaner’s prediciton model. We plot the average movements for each feature and for each dataset in Figure 2. For the Wine quality, we see the most movement in alcohol followed closely by volatile acidity and residual sugar. These results give the insight into how much an adversary must modify each feature in order to fool the learner. It is clear, for example, that the alcohol content must be modified considerably more than the citric acid to generate the most effective wine. On the other hand, the citric acid content required will little change to achieve a value sufficient to fool the learner’s predictor. The citric acid, therefore, might be considered a particularly vulnerable feature that could easily be exploited and perhaps should be treated with more scrutiny. In the real estate data, we see the distance to the neaest MRT station as the most resilient feature, requiring relatively substantial change to effect the learner’s predictor. Meanwhile, the transaction data and longitude appear to be the easiest to exploit.
Figure 2: Feature modifications made by the adversary to trick the learner
It is clear from the experimental results that the constrained pessimistic bilevel model provides an effective method of accounting for adversarial manipulation during the training process of regression scenarios. The lower-level successfully anticipates the movements of an adversary, which leads to improved performance by the prediction model trained in the upper-level. We see a significant change in performance as we increase the number of instances available to be manipulated by the adversary. In particular, we note that a balance needs to be struck between allowing the adversary enough instances that they have sufficient influence over then training process, while not too many that the model overestimates the influence of the adversary. Finally, we demonstrated how, unlike the work in [37], we could investigate the values of the adversary’s data at the solution. This allowed us to identify which features were most vulnerable to exploitation.
4 Conclusion
Throughout this paper, we demonstrated how the pessimistic bilevel model proposed by [37] can be adapted to adversarial regression scenarios. In particular, by anticipating adversarial movement in the lower-level, we can train prediction functions that are resilient to adversaries who attempt to influence particular outcomes from a prediction model by strategically transforming their data. For example, in the context of property valuation, an adversary might purposefully provide falsified information to be achieve higher valuations and hence charge a higher selling price. While there exist pessimistic bilevel approaches to adversarial evasion scenarios, these models exploit strong assumptions about the convexity of the adversary’s problem and the uniqueness of their solution to reformulate the program into its optimistic variant, which, while easier to solve, oversimplifies the capabilities of the adversary. Since our model makes no such assumptions, we retain the pessimistic aspect of the bilevel model, allowing us to accurately capture the antagonistic nature of these scenarios.
To assess the again simulated adversarial influence on a number of applications, including insurance appraisal, wine quality appraisal and real estate pricing. Throughout numerical experiments, we demonstrated the ability of this model to train resilient predictors that provided improved performance over traditionally-trained predictors. Further to this, we again investigated varying the number of samples available to the adversary and the similarity threshold and observed a clear pattern in the form of a trade-off between allowing the adversary enough freedom and influence to generate sufficient movement that impacts the training process, while not too much such that the adversary provides too much influence over the training process, leading to overly pessimistic predictors.
Appendix A Derivatives for the leader and follower
Let $x\in\mathbb{R}^{q}$ be a sample of data with corresponding label $y\in\mathbb{R}$ abd let $w\in\mathbb{R}^{q}$ be the weights of the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
$\sigma(w,x)=w^{T}x.$
We define the upper-level (learner’s) loss function $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
$\mathcal{L}(\sigma(w,x),y):=\|\sigma(w,x)-y\|_{2}^{2}.$
Let $z\in\mathbb{R}$ be the adversary’s target label. We define the lower-level (adversary’s) loss function $\ell:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
$\ell(\sigma(w,x),y):=\|\sigma(w^{T}x)-z\|_{2}^{2}.$
Let $D\in\mathbb{R}^{mq}$ be the static data with corresponding labels $\gamma\in\mathbb{R}^{n}$ and let $X\in\mathbb{R}^{mq}$ be the adversary’s data with corresponding labels $Y\in\mathbb{R}^{m}$. The upper-level objective $F:\mathbb{R}^{q}\times\mathbb{R}^{mq}\rightarrow\mathbb{R}$ is defined as,
$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i})+\frac{1}{\rho}\|w\|_{2}^{2}.$
The derivative of the upper-level (leader’s) loss with respect to the learner’s weights is given by,
$\frac{\partial F}{\partial w}(w,X)=\frac{2}{n}D^{T}(w^{T}D-\gamma)+\frac{2}{m}X^{T}(w^{T}X-Y)+\frac{2}{\rho}w.$
The second derivative of the upper-level loss with respect to the learner’s weights is given by
$\frac{\partial^{2}F}{\partial ww}(w,X)=\frac{2}{n}D^{T}D+\frac{2}{m}X^{T}X+\frac{2}{\rho}$
The derivative of the upper-level loss function with respect to the adversary’s data is given by
$\frac{\partial\mathcal{L}(w,X)}{\partial X_{ij}}=\frac{2}{m}w_{j}(w^{T}X_{i}-y_{i}),\quad i=1,\dots,m,\;j=1,\dots,q.$
The second derivative with respect to the adversary’s data is given by the following cases,
$\frac{\partial^{2}\mathcal{L}}{\partial X_{ij}X_{kl}}=\begin{cases}\frac{2}{m}w_{j}w_{l}&i=k\\ 0&\text{otherwise}\end{cases}.$
Let $X^{0}\in\mathbb{R}^{mq}$ be the start point of the adversary’s data. We define the constraints $g:\mathbb{R}^{mq}\rightarrow(-1,1)$ of the lower-level component-wise as
$g_{i}(X)=\delta-d(X_{i},X_{i}^{0}),\;i\in\{1,\dots,m\},$
where
$d(X_{i},X_{i}^{0})=\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i\in\{1,\dots,m\},$
and $\delta\in\mathbb{R}$ is the similarity threshold. The derivative of the constraints with respect to the classifier weights is 0. The derivative with respect to the adversary’s data is given by the following cases,
$\frac{\partial g_{i}(X)}{\partial X_{jk}}=\begin{cases}\frac{X_{ik}^{0}}{\|X_{i}\|\cdot\|X_{i}^{0}\|}-d(X_{i},X_{i}^{0})\frac{X_{ik}}{\|X_{i}\|^{2}}&i=j\\ 0&i\neq j\end{cases}.$
The second derivative with respect to the adversary’s data is given by the cases,
$\frac{\partial g_{i}(X)}{X_{jk}X_{lc}}=\begin{cases}\frac{X_{ic}X_{ik}^{0}+X_{ik}X_{ic}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}X_{ic}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}&i=j=l,k\neq c\\ \frac{2X_{ik}X_{ik}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}^{2}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}+\frac{d(X_{i},X_{i}^{0})}{\|X_{i}\|^{2}}&i=j=l,k=c\\ 0&\text{otherwise}.\end{cases}$